media%2Fe83%2Fe8353209-1dc8-4794-9ade-69

media%2Ff7e%2Ff7eadf98-aa1d-447f-81ad-21
Answer #1

Given,

Permeate flow rate = 12 MGD

Feed Salt concentration = 2600 mg/L

Plant recovery (r) = 84%

Average Salt Removal (R) = 93 %

29) Average Salinity of the total permeate from the plant:

= Feed Salt concentration * (1-R) = 2600 * (1 - 0.93) = 182 mg/L

30) Average Salinity of the total concentrate from the plant:

= Feed Salt concentration / (1-r) = 2600/(1 - 0.84) = 16,250 ~ 16,500 mg/L

31) Total mass rate of salt in concentrate stream:

= (Feed Salt concentration * Feed Flow rate) - (Permeate flow rate * Permeate flow rate) = (0.021698 lb/gal * 14.3 * 106 gal/day) - (0.001519 lb/gal * 12 * 106 gal/day) = (310281.4) - 18228 = 292,053 lb/day ~ 291,700 lb/day

32) Feed Flow rate:

= Permeate flow rate/r = 12/0.84 = 14.3 MGD

33) Concentrate Flow rate:

= Feed flow rate - Permeate flow rate = 14.3 - 12 = 2.3 MGD